Solve for x
log x +log(x-5)=log14

Answer by ewatrrr(24785)


Hi
Solve for x
log x +log(x-5)=log14
log x(x-5)/14 = 0
10^0 = x(x-5)/14
1 = x(x-5)/14
14 = x(x-5)
x^2 - 5x - 14 = 0
factoring
(x+2)(x-7)=0 Note:SUM of the inner product(2x) and the outer product(-7x) = -5x
(x+2)= 0 negative solution is extraneous
(x-7)=0
x = 7